OpenStudy (anonymous):

How to convert a Standard Equation to Slope Interecpt Form? Someone help please!?

OpenStudy (anonymous):

Ax +By = C Divide both sides by B A/Bx +y = C/B Now put eq in y = mx+b form y = -A/B x +C/B

OpenStudy (anonymous):

Hmm

OpenStudy (anonymous):

4x + y = 8

OpenStudy (anonymous):

How would I do this one, since Y is on it's own?

jimthompson5910 (jim_thompson5910):

solve for y and get the equation into y = mx+b form 4x + y = 8 4x + y-4x = 8-4x ... Subtract 4x from both sides. y = -4x + 8

jimthompson5910 (jim_thompson5910):

so 4x + y = 8 is equivalent to y = -4x+8 y = -4x+8 matches up with y = mx+b where m = -4 (slope) b = 8 (y-intercept)